OpenStudy (anonymous):

Find the simplest form of this expression. sqrt2*sqrt9

OpenStudy (bahrom7893):

sqrt(2*9)=3*sqrt(2*3)=3sqrt(6)

OpenStudy (id21):

\[\sqrt{2}\times \sqrt{9}=3\sqrt{2}\]